# On Positional Bias of Faithfulness for Long-form Summarization

### v1

@@ A.2 Model Details
For all models, we use the default generation methods. For open-source models, we use the available Huggingface repository for Llama-3.1-8B [^3] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct and 70B [^4] https://huggingface.co/meta-llama/Llama-3.1-70B-Instruct, and also for Mixtral [^5] https://huggingface.co/mistralai/Mixtral-8x7B-Instruct-v0.1 and Mixtral Large. [^6] https://huggingface.co/mistralai/Mixtral-8x22B-Instruct-v0.1 For GPT-4o, we use gpt-4o as of October 13th, 2024. We run with bfloat16 and use 8 A100s to run all generations and evaluations.
@@ Appendix B Additional Metric Results
MiniCheck Original Min 41.1 56.7 66.9 72.2 68.0 77.5 50.5 46.2 59.9
Original Mean 52.5 62.8 60.8 72.2 56.1 77.5 49.5 48.6 60.0
Original Max 58.9 63.3 50.0 72.2 50.0 77.5 50.0 50.6 59.0
Min Min 49.4 56.7 66.9 50.8 50.0 77.5 52.8 50.3 56.8
Min Mean 50.1 62.8 60.8 50.8 50.0 77.5 54.2 49.6 57.0
Min Max 59.7 63.3 50.0 50.8 53.0 77.5 54.2 50.6 57.4
Mean Min 49.4 56.7 66.9 51.5 50.0 77.5 52.8 48.6 56.7
Mean Mean 54.8 62.8 60.8 51.5 55.5 77.5 54.2 49.4 58.3
Mean Max 64.3 63.3 50.0 51.5 55.5 77.5 64.8 55.2 60.3
Max Min 38.7 56.7 66.9 73.1 66.0 77.5 50.9 46.6 59.5
Max Mean 54.7 62.8 60.8 73.1 62.3 77.5 56.9 43.5 61.4
Max Max 53.2 63.3 50.0 73.1 50.0 77.5 50.0 50.6 58.4
@@ B.1 Results on Chunking
B.2 Full results
We present the full results, including an exploration of summary sentence-level merging strategies, in Table 7. In the case of MiniCheck, taking the mean over the summary sentences yields the highest correlations in two out of four cases, while taking the maximum does so in the remaining two cases. For GPT-4o, we observe that, across all document merging strategies, averaging over the summary sentences achieves the highest correlations overall. This finding is consistent with prior work.
@@ Appendix C Additional Faithfulness Analysis Results
C.1 Faithfulness Metric as Alignment and Attribution
@@
Figure 6: Coverage analysis across different positions.
C.2 Coverage Analysis
@@
C.3 Faithfulness Analysis using GPT-4o

### v2

@@ 7 Conclusion
Acknowledgement
We thank the anonymous reviewers for their valuable feedback. We also would like to thank Philippe Laban and Kung-Hsiang Huang for the helpful discussions and for providing help with the datasets SummHay and DiverseSumm, respectively.
@@ A.2 Model Details
For all models, we use the default generation methods. For open-source models, we use the available Huggingface repository for Llama-3.1-8B [^4] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct and 70B [^5] https://huggingface.co/meta-llama/Llama-3.1-70B-Instruct, and also for Mixtral [^6] https://huggingface.co/mistralai/Mixtral-8x7B-Instruct-v0.1 and Mixtral Large. [^7] https://huggingface.co/mistralai/Mixtral-8x22B-Instruct-v0.1 For GPT-4o, we use gpt-4o as of October 13th, 2024. We run with bfloat16 and use 8 A100s to run all generations and evaluations. The approximate costs for GPT-4o are as follows: metric full setting costs \$37.1, metric splitting the document setting costs \$79.9, generation costs \$ 15.2, and generation with different methods costs \$97.2.
@@ Appendix B Additional Metric Results
MiniCheck Original Min 47.8 56.6 64.7 79.7 76.0 83.0 46.8 53.5 63.5
Original Mean 56.1 55.7 55.1 79.7 57.0 83.0 57.9 54.1 62.3
Original Max 53.8 53.9 50.0 79.7 50.0 83.0 50.9 55.2 59.6
Min Min 49.4 56.6 64.7 50.0 50.0 83.0 51.4 49.8 56.9
Min Mean 54.7 55.7 55.1 50.0 50.0 83.0 54.2 49.6 56.5
Min Max 60.3 53.9 50.0 50.0 50.8 83.0 54.2 53.8 57.0
Mean Min 49.4 56.6 64.7 52.6 50.0 83.0 51.4 49.3 57.1
Mean Mean 64.5 55.7 55.1 52.6 53.4 83.0 54.2 46.0 58.1
Mean Max 60.4 53.9 50.0 52.6 46.2 83.0 66.7 51.2 58.0
Max Min 39.4 56.6 64.7 84.8 69.7 83.0 53.2 47.1 62.3
Max Mean 56.6 55.7 55.1 84.8 62.3 83.0 59.7 49.3 63.3
Max Max 53.2 53.9 50.0 84.8 50.0 83.0 50.9 54.2 60.0
@@ B.1 Results on Chunking
B.2 Full Metric Results
We present the full results, including an exploration of summary sentence-level merging strategies, in Table 8. For MiniCheck, the original chunking method using the minimum function over summary sentences yields the highest correlation—though this is largely driven by high accuracy on ArXiv. Meanwhile, the maximum strategy with mean aggregation, which previous studies have identified as most effective, falls short by only 0.2 points. For GPT-4o, we observe that, across all document merging strategies, averaging over the summary sentences achieves the highest correlations overall. This finding is consistent with prior work.
@@ B.2 Full Metric Results
Figure 5: Faithfulness analysis across different positions of documents with chunking.
@@ Appendix C Additional Faithfulness Analysis Results
Task R1 R2 RL
MultiXScience 24 24 12
PubMed 26 26 13
MultiNews 18 28 13
ArXiv 32 32 13
DiverseSumm 41 41 21
Table 9: Similarity of the documents using ROUGE.
C.1 Content Overlap Analysis
Taking the maximum faithfulness implicitly assumes that the conent do not overlap. To verify the variability in similarity, we calculate ROUGE-1/2/L between all document pairs and aggregate the results to show the variablity for each task . We exclude SummHay as it is prohibitively expensive to calculate this for 100 documents. Table 9 shows the results. Examining unigrams (R1), we observe a high degree of overlap, particularly for DiverseSumm. This is expected, as all 10 documents focus on the same news event. We note that while such overlap may not be covered by only considering the highest faithfulness for one of the documents, we evaluate faithfulness relative to a single document to minimize the influence of coverage, as mentioned in Section 4.2. Determining how many documents a summary can be attributed to is complex and would require defining a threshold, which we leave for future work. Nevertheless, the coverage analysis shown in Figure 7 can be considered the case where the summary is related to all documents.
C.2 Faithfulness Analysis with Chunking
We perform the analysis with calculating the faithfulness with chunking, instead of natural boundaries. We show it in Figure 5.
@@ C.2 Faithfulness Analysis with Chunking
Figure 7: Coverage analysis across different positions.
C.3 Discussions on Variance Across Tasks
In Figure 2, we observe a large variance across the different tasks. We provide a discussion on possible reasons.
Context Length.
Based on the dataset statistics provided earlier, we analyzed the tasks from those with the shortest document lengths to those with the longest. This reveals an intriguing trend: MultiXScience, shown in Figure 2, exhibits primarily a lead bias due to its short input length. As we move to datasets with longer input lengths, the U-shaped trend becomes more apparent. MultiXScience can thus be interpreted as representing the early stage of this trend, which evolves as context length increases. We demonstrate this similar progression in Section 4.4, where increasing contexts shift the observed behavior. The downward trend of MultiXScience in Figure 3 can similarly be attributed to insufficient context length, leaving it in the early lead bias phase.
Task Type.
Another observable trend is the variation across different types of summarization tasks. For example, ArXiv and PubMed involve single long-document summarization, whereas MultiNews, MultiXScience, DiverseSumm, and SummHay are classified as MDS. Our findings indicate that single-document summarization tasks exhibit more consistent model behavior and lower variance, while MDS tasks show much higher variance. We hypothesize that this increased variability stems from the need for synthesis across multiple documents, which impacts the faithfulness of generated summaries.
C.4 Faithfulness Metric as Alignment and Attribution
@@
C.5 Coverage Analysis
@@
C.6 Faithfulness Analysis using GPT-4o